# X-ray Dips and Polarization Angle Swings in GX 131

### v1

@@ 1 Introduction
Low Mass X-ray Binaries (LMXBs) are systems in which a compact object accretes matter from a low mass companion star via Roche-lobe overflow and a disk. Systems hosting a weakly magnetized neutron star (WMNS) comprise two main classes, named Z and Atoll sources, after the characteristic shape they trace in the X-ray color-color diagram (CCD). Z sources are persistent and bright, displaying luminosities close to the Eddington limit, whereas Atoll sources are about a decade less luminous, frequently display thermonuclear bursting activity and in some cases undergo transient outbursts; see, e.g., van der Klis (1989). The millisecond spin periods that are expected to characterize WMNS in LMXBs have been detected only from a few dozen Atoll sources through coherent pulsations during transient outbursts and/or burst oscillations during thermonuclear bursts (Papitto et al., 2014). Fast aperiodic and quasi-periodic variability is detected in the range between few Hz and kHz and presents clear similarities in both Z and Atoll sources. The energy spectra of accreting WMNSs can be described by a model consisting of a multicolor blackbody, representing the thermal emissions from the inner accretion disk regions, and a higher-energy Comptonized continuum, which dominates the spectrum and likely originates from the boundary layer between the disk and the NS (BL; Shakura & Sunyaev, 1988; Popham & Sunyaev, 2001) or the plasma layer at the NS surface, the so-called spreading layer (SL; Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006). Reflection spectrum (White et al., 1988; George & Fabian, 1991; Fabian et al., 1989; Giridharan et al., 2023; Ludlam, 2024) caused by radiation intercepted and reprocessed by the accretion disk is also observed in many cases. The presence of powerful outflows, in the form of collimated fast jets emitted perpendicularly to the disk (Baglio et al., 2017) and/or fan-shaped slower winds launched from moderate altitudes above the disk (Allen et al., 2018) have also been established in a number of BH and NS LMXBs.
@@
X-ray polarimetry has introduced a novel diagnostic for NS LMXBs through two new observables, the polarization angle (PA) and the polarization degree (PD), and their variability. This new information can provide insight into the accretion physics of these objects and the geometry of the BL (and/or SL) and possibly of the ADC and the disk wind. IXPE has already observed several systems. For Cyg X-2, the polarization angle (PA) was aligned along the position angle of the radio jet (Farinelli et al., 2023), which is likely perpendicular to the accretion disk. On the other hand, La Monaca et al. (2024) observed a significant difference between the PA measured from Sco X-1 and the jet position angle, suggesting a more complex geometry for this source. In Cir X-1 (Rankin et al., 2024), XTE J1701462 (Cocchi et al., 2023) and GX 51 (Fabiani et al., 2024) the polarization varied with the state and/or the hardness of the source with a PA rotation up to 67° in Cir X-1 and PD passing from in HB to in NB/FB of Z sources (Di Marco & IXPE Science Team, 2024). In the Atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820303 (Di Marco et al., 2023a), and 4U 162449 (Saade et al., 2024), and the Z source GX 340+0 (La Monaca, Fabio et al., 2024) a dependence of the PD on energy was observed.
@@ 2 Observations and data reduction
The IXPE payload consists of three identical telescopes, providing imaging and spectral polarimetry over the 2–8 keV energy band (Weisskopf et al., 2022; Soffitta et al., 2021). It observed GX 131 for the third time (Observation ID 03003401) in the period 2024 April 20 20:49 UTC – April 23 01:01 UTC (orbital phase 0.33–0.42), for a total effective exposure time of 100 ks for each telescope (see Table 1). The 2–8 keV IXPE light curve binned every 200 s is shown in Figure 1, along with the hardness ratio of the rates in the 2–4 and 4–12 keV bands.
@@
IXPE data were processed for polarimetric analysis using the ixpeobssim package version 31.0.1 (Baldini et al., 2022). For spectral and spectropolarimetric analysis, data extraction was performed using HEASoft version 6.33.2 and the standard ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc). In the case of IXPE, we used the CALDB released on 2024 February 28. Source events were extracted from IXPE data by selecting a circular region of radius 100″ centered on the position of the source, using SAOImageDS9.222https://sites.google.com/cfa.harvard.edu/saoimageds9 Due to the high flux of the source, the background was not subtracted, as prescribed by Di Marco et al. (2023b). In the model-independent polarimetric analysis, based on the pcube algorithm in ixpeobssim, the unweighted analysis was adopted, whereas for the model-dependent spectral and spectropolarimetric analyses we used the ‘weighted analysis’. The latter is based on the ellipticity of the photoelectrons tracks collected by the IXPE DUs and follows the approach of Di Marco et al. (2022). Data were binned so as to have at least 30 counts per energy channel in the spectral analysis, while in the spectropolarimetric analysis we applied a constant energy binning of 120 eV for all three energy distributions of the Stokes parameters , , and , given the lower statistic available in the and spectra.
The extraction of source and background events was performed in the Swift–XRT data with SAOImageDS9 using an annulus (inner radius of about 5″ and an outer radius of about 40″) in order to avoid possible pile-up effects (Romano et al., 2006); the source annulus was centered at the GX 131 coordinates, the background annulus about 80″ off-source. The Swift–XRT data were fitted in the 0.8–8 keV band to optimize the statistics and grouped so as to have at least 30 counts per bin. The response matrices released in the HEASARC CALDB on 2023 July 25 were applied in the analysis.
@@ 3.2 Spectral analysis
The NICER data partially covered the IXPE observation with several snapshots, some of which were during dips (see Table 1 and Figure 1). To model the spectra in the Dip and Off-dip states, the data were selected on the basis of their HR, as shown in Figure 1. By virtue of its better spectral capabilities and throughput (compared to IXPE and Swift–XRT) NICER resolved some of the absorption features reported in the literature (Díaz Trigo et al., 2012; D’Aì et al., 2014; Saavedra et al., 2023). In particular, residuals in the 6-8 keV region were clearly visible when the continuum model was fitted to the spectra. To account for those features, we followed the approach in Díaz Trigo et al. (2012) and included a warmabs component to model absorption by photoionized plasma along the line of sight. Moreover, a cabs component was included to account for the effects of Thomson/Compton scattering, which was found significant when modeling both the persistent and dipping emission (Boirin et al., 2005; Díaz Trigo et al., 2006, 2012). The latter component was forced to have the column density of the warmabs component (which accounts for the number of electrons per hydrogen atom for solar abundances material Stelzer et al., 1999). Furthermore, a gaussian component was included to account for a broad Fe line emission. The complete model we fit to the NICER spectra was thus TBabs*warmabs*cabs*(diskbb+bbodyrad+gauss). The best-fit results reported in Table 3 are in general agreement with those from XMM–Newton and Chandra observations (Díaz Trigo et al., 2012; D’Aì et al., 2014). The spectra are shown in Figure 6. Absorption was higher during the dips, especially during the first one, whereas the diskbb temperature was lower, keV, than outside the dips ( keV). The bbodyrad component remained instead at a nearly constant temperature of keV.
@@ 3.3 Spectropolarimetric analysis
As a second step, we associated separate polconst polarization components to the diskbb and bbodyrad components: tbabs*(polconst*diskbb+polconst*bbodyrad+gauss). The best-fit results for this model are reported in Table 4. Polarization was detected at a level of across the Off-dip and Dip states in the harder component (bbodyrad), together with marginal evidence for a PA change of . For the diskbb component we found in the Dip state, whereas in the Off-dip state we derived an upper limit of at 90% CL. We warn that the soft component’s 10% PD during the dip should be considered with care, as the polarizations of the two components in the fit result are almost orthogonal, meaning that they subtract from each other and partial degeneracy sets in.

### v2

@@ 1 Introduction
Low Mass X-ray Binaries (LMXBs) are systems in which a compact object accretes matter from a low mass companion star via Roche-lobe overflow and a disk. Systems hosting a weakly magnetized neutron star (WMNS) comprise two main classes, named Z and Atoll sources, after the characteristic shape they trace in the X-ray color-color diagram (CCD). Z sources are persistent and bright, displaying luminosities close to the Eddington limit, whereas Atoll sources are about a decade less luminous, frequently display thermonuclear bursting activity and in some cases undergo transient outbursts; see, e.g., van der Klis (1989). The millisecond spin periods that are expected to characterize WMNSs in LMXBs have been detected only from a few dozen Atoll sources through coherent pulsations during transient outbursts and/or burst oscillations during thermonuclear bursts (Papitto et al., 2014). Fast aperiodic and quasi-periodic variability is detected in the range between few Hz and kHz and presents clear similarities in both Z and Atoll sources. The energy spectra of accreting WMNSs can be described by a model consisting of a multicolor blackbody, representing the thermal emissions from the inner accretion disk regions, and a higher-energy Comptonized continuum, which dominates the spectrum and likely originates from the boundary layer between the disk and the NS (BL; Shakura & Sunyaev, 1988; Popham & Sunyaev, 2001) or the plasma layer at the NS surface, the so-called spreading layer (SL; Inogamov & Sunyaev, 1999; Suleimanov & Poutanen, 2006). Reflection spectrum (White et al., 1988; George & Fabian, 1991; Fabian et al., 1989; Giridharan et al., 2023; Ludlam, 2024) caused by radiation intercepted and reprocessed by the accretion disk is also observed in many cases. The presence of powerful outflows, in the form of collimated fast jets emitted perpendicularly to the disk (Baglio et al., 2017) and/or fan-shaped slower winds launched from moderate altitudes above the disk (Allen et al., 2018) have also been established in a number of BH and NS LMXBs.
@@
X-ray polarimetry has introduced a novel diagnostic for NS LMXBs through two new observables, the polarization angle (PA) and the polarization degree (PD), and their variability. This new information can provide insight into the accretion physics of these objects and the geometry of the BL (and/or SL) and possibly of the ADC and the disk wind. IXPE has already observed several systems. For Cyg X-2, the polarization angle (PA) was aligned along the position angle of the radio jet (Farinelli et al., 2023), which is likely perpendicular to the accretion disk. On the other hand, La Monaca et al. (2024) observed a significant difference between the PA measured from Sco X-1 and the jet position angle, suggesting a more complex geometry for this source. In Cir X-1 (Rankin et al., 2024), XTE J1701462 (Cocchi et al., 2023) and GX 51 (Fabiani et al., 2024) the polarization varied with the state and/or the hardness of the source with a PA rotation up to 67° in Cir X-1 and PD passing from in the horizontal branch to in the Normal/Flaring branch of Z sources (Di Marco & IXPE Science Team, 2024). In the Atoll sources GX 9+9 (Ursini et al., 2023), 4U 1820303 (Di Marco et al., 2023a), and 4U 162449 (Saade et al., 2024), and the Z source GX 340+0 (La Monaca, Fabio et al., 2024) a dependence of the PD on energy was observed.
@@ 2 Observations and data reduction
The IXPE payload consists of three identical telescopes, providing imaging and spectral polarimetry over the 2–8 keV energy band (Weisskopf et al., 2022; Soffitta et al., 2021). It observed GX 131 for the third time (Observation ID 03003401) in the period 2024 April 20 20:49 UTC – April 23 01:01 UTC (orbital phase 0.33–0.42), for a total effective exposure time of 100 ks for each telescope (see Table 1). The 2–8 keV IXPE light curve binned every 200 s is shown in Figure 1, along with the hardness ratio of the rates in the 2–4 and 4–8 keV bands.
@@
IXPE data were processed for polarimetric analysis using the ixpeobssim package version 31.0.1 (Baldini et al., 2022). For spectral and spectropolarimetric analysis, data extraction was performed using HEASoft version 6.33.2 and the standard ftools (Nasa High Energy Astrophysics Science Archive Research Center (2014), Heasarc). In the case of IXPE, we used the CALDB released on 2024 February 28. Source events were extracted from IXPE data by selecting a circular region of radius 100″ centered on the position of the source, using SAOImageDS9.222https://sites.google.com/cfa.harvard.edu/saoimageds9 Due to the high flux of the source, the background was not subtracted, as prescribed by Di Marco et al. (2023b). In the model-independent polarimetric analysis, based on the pcube algorithm in ixpeobssim, the unweighted analysis was adopted, whereas for the model-dependent spectral and spectropolarimetric analyses we used the ‘weighted analysis’. The latter is based on the ellipticity of the photoelectrons tracks collected by the IXPE detector units (DUs) and follows the approach of Di Marco et al. (2022). Data were binned so as to have at least 30 counts per energy channel in the spectral analysis, while in the spectropolarimetric analysis we applied a constant energy binning of 120 eV for all three energy distributions of the Stokes parameters , , and , given the lower statistic available in the and spectra.
The extraction of source and background events was performed in the Swift–XRT data with SAOImageDS9 using an annulus (inner radius of about 5″ and an outer radius of about 40″) in order to avoid possible pile-up effects (Romano et al., 2006); the source annulus was centered at the GX 131 coordinates, and the background annulus about 80″ off-source. The Swift–XRT data were fitted in the 0.8–8 keV band to optimize the statistics and grouped so as to have at least 30 counts per bin. The response matrices released in the HEASARC CALDB on 2023 July 25 were applied in the analysis.
@@ 3.2 Spectral analysis
The NICER data partially covered the IXPE observation with several snapshots, some of which were during dips (see Table 1 and Figure 1). To model the spectra in the Dip and Off-dip states, the data were selected on the basis of their hardness ratio (HR), as shown in Figure 1. By virtue of its better spectral capabilities and throughput (compared to IXPE and Swift–XRT) NICER resolved some of the absorption features reported in the literature (Díaz Trigo et al., 2012; D’Aì et al., 2014; Saavedra et al., 2023). In particular, residuals in the 6-8 keV region were clearly visible when the continuum model was fitted to the spectra. To account for those features, we followed the approach in Díaz Trigo et al. (2012) and included a warmabs component to model absorption by photoionized plasma along the line of sight. Moreover, a cabs component was included to account for the effects of Thomson/Compton scattering, which was found to be significant when modeling both the persistent and dipping emission (Boirin et al., 2005; Díaz Trigo et al., 2006, 2012). The latter component was forced to have the column density of the warmabs component (which accounts for the number of electrons per hydrogen atom for solar abundances material Stelzer et al., 1999). Furthermore, a gaussian component was included to account for a broad Fe line emission. The complete model we fit to the NICER spectra was thus TBabs*warmabs*cabs*(diskbb+bbodyrad+gauss). The best-fit results reported in Table 3 are in general agreement with those from XMM–Newton and Chandra observations (Díaz Trigo et al., 2012; D’Aì et al., 2014). The spectra are shown in Figure 6. Absorption was higher during the dips, especially during the first one, whereas the diskbb temperature was lower, keV, than outside the dips ( keV). The bbodyrad component remained instead at a nearly constant temperature of keV.
@@ 3.3 Spectropolarimetric analysis
As a second step, we associated separate polconst polarization components to the diskbb and bbodyrad components: tbabs*(polconst*diskbb+polconst*
bbodyrad+gauss).
The best-fit results for this model are reported in Table 4. Polarization was detected at a level of across the Off-dip and Dip states in the harder component (bbodyrad), together with marginal evidence for a PA change of . For the diskbb component we found in the Dip state, whereas in the Off-dip state we derived an upper limit of at 90% CL. We warn that the soft component’s 10% PD during the dip should be considered with care, as the polarizations of the two components in the fit result are almost orthogonal, meaning that they subtract from each other and partial degeneracy sets in.